1. When H2(g) reacts with Cl2(g) to form HCl(g) , 185 kJ of energy are evolved for each mole of H2(g) that reacts. Write a balanced thermochemical equation for the reaction with an energy term in kJ as part of the equation. Note that the answer box for the energy term is case sensitive.
Use the SMALLEST INTEGER coefficients possible and put the energy term (including the units) in the last box on the appropriate side of the equation. If a box is not needed, leave it blank.
2. When CO(g) reacts with H2O(l) to form CO2(g) and H2(g), 2.80 kJ of energy are absorbed for each mole of CO(g) that reacts. Write a balanced thermochemical equation for the reaction with an energy term in kJ as part of the equation. Note that the answer box for the energy term is case sensitive.
1
Expert's answer
2019-10-31T07:29:39-0400

1) H2(g)+Cl2(g)=2HCl+ 185kJ

2) CO(g)+H2O(l)=CO2(g)+H2(g)- 2.80 kJ